OpenStudy (nottim):

Find the object's speeds v1, v2, and v3 at times t1=2.0s, t2=4.0s, and t3=13s.

OpenStudy (anonymous):

v=x/t so to find v from a x-t graph u need to find the slope of the graph at that point. for example,at t1=2s ---> u see that the graph has the same slant(slope=change in y/change in x), so in the interval in which graph has that constant slope pick any two points and apply the formula of slope to find v i choose points (7,3) and (2,1),so slope(v)=(3-1)/(7-2)=2/5=0.4 m/s